# Shaping the Safety Boundaries: Understanding and Defending Against Jailbreaks in Large Language Models

### v1

@@ Appendix B Full View of Activation Space
Appendix C Supplementary Experiments
@@ C.2 Inclusion Ratio Experiments
For a layer $l$, to measure the portion of a set of jailbreak activations $A^{l}=\{a^{l}_{0},a^{l}_{1},\cdots,a^{l}_{n}\}$ that resides in harmful activation space, we propose an inclusion ratio. Based on 8,556 harmful samples gathered in Table 3, we calculate a ball that covers $80\%$ activations. The center of the ball is $\mu_{\mathcal{D}}^{l}$, and the radius of the ball is denoted as $r_{\mathcal{D}}^{l}$.
@@
$\rho^{l}_{\mathrm{inclusion}}=\frac{\left|\{a^{l}_{i}\in A^{l}\mid\|a^{l}_{i}-\mu_{\mathcal{D}}^{l}\|_{2}\leq r_{\mathcal{D}}^{l}\}\right|}{|A^{l}|},$
where $\|a^{l}_{i}-\mu_{\mathcal{D}}^{l}\|_{2}$ is the distance between the activation and the center. We calculate inclusion ratios of different types of jailbreak activations, with three representative types shown in Figure 1. We then apply ABD on all types of jailbreak activations and calculate their inclusion ratios. Notably, despite different types of jailbreaks, they all achieve an inclusion ratio of $100\%$ with ABD applied, which verifies the effect of our method.
Appendix D Supplementary Illustration on ABD
@@
D.1 Visualization of the penalty function.
@@
D.2 ABD Optimization Settings
@@
We adopt Optuna Akiba et al. (2019) as our framework of optimization. We follow Optuna’s default settings, i.e., Gaussian process-based Bayesian Optimization.
Appendix E Supplementary Illustration on Experiments
E.1 Generation Configs
When conducting experiments, we directly utilize most of the original configurations of Vicuna-7B-v1.3 and Llama2-7B-chat. Specifically, we set max_new_tokens=128. We find that the proper functioning of these LLMs largely depends on the chat template of the inputs. We apply chat templates in fastchat v0.2.36 by:
@@ E.1 Generation Configs
We set template_name="vicuna" for Vicuna-7B-v1.3 and template_name="llama-2" for Llama2-7B-chat.
E.2 Jailbreak Methods
@@ E.3 Defense Methods
E.4 Metrics
E.4.1 Calculating DSR
@@
We adopt refusal strings as proposed by Xu et al. (2024), presented in Table
E.4.2 Calculating Overall Scores
@@
where $\mathrm{t}$ is Runtime per Query and $\mathrm{Avg.}$ represents overall averaged Just-Eval scores. 4.

### v2

@@ Appendix B Full View of Activation Space
Appendix C Supplementary Observational Experiments
@@ C.2 Inclusion Ratio Experiments
For a layer $l$, to measure the portion of a set of jailbreak activations $A^{l}=\{a^{l}_{0},a^{l}_{1},\cdots,a^{l}_{n}\}$ that resides in harmful activation space, we propose an inclusion ratio. Based on 8,556 harmful samples gathered in Table 4, we calculate a ball that covers $80\%$ activations. The center of the ball is $c_{\mathcal{D}}^{l}$, and the radius of the ball is denoted as $r_{\mathcal{D}}^{l}$.
@@
$\rho^{l}_{\mathrm{inclusion}}=\frac{\left|\{a^{l}_{i}\in A^{l}\mid\|a^{l}_{i}-c_{\mathcal{D}}^{l}\|_{2}\leq r_{\mathcal{D}}^{l}\}\right|}{|A^{l}|},$
where $\|a^{l}_{i}-c_{\mathcal{D}}^{l}\|_{2}$ is the distance between the activation and the center. We calculate inclusion ratios of different types of jailbreak activations, with three representative types shown in Figure 1. We then apply ABD on all types of jailbreak activations and calculate their inclusion ratios. Notably, despite different types of jailbreaks, they all achieve an inclusion ratio of $100\%$ with ABD applied, which verifies the effect of our method.
C.3 Ablation on Optimizing Objective
We conduct ablation studies to determine the best $w$ in $\mathcal{J}_{\mathrm{total}}(\Theta,M)$. We adopt Vicuna-7B-v1.3 as the backbone model, and use the 50-behavior subset as the testing set. Additionally, we fix optimization steps to 2000. DSR results are shown in Table 5. We discover that (1) Higher w improves robustness on attacked datasets. (2) Lower w preserves performance on unattacked datasets. (3) At w=0.8, the model achieves the best average DSR, balancing robustness and benign performance. Therefore, we set $w=0.8$ in our experiments in Section 5.
Jailbreak No defense $w=0.2$ $w=0.5$ $w=0.8$ $w=1$
No Jailbreaks 96% 100% 86% 100% 86%
GCG-Individual 7% 75% 88% 93% 100%
Avg. 52% 88% 87% 96% 93%
Table 5: Defense performance comparison under different $w$.
Appendix D Supplementary Illustration on Safety Boundaries
To gain a deeper understanding of the safety boundary, we present a case study indicating how the model’s responses vary based on the level of RAS applied to layer 1. The results are shown in Table 6. It shows that (1) The safety boundary is implicit, as a significant shift does not necessarily results in successful deception, such as when $r=80$. (2) As the responses move further away from this boundary, the model begins to exhibit uncommon response patterns such as giving a presentation ($r=100$) or more casual styles ($r=30,r=80$). We believe this phenomenon may correlate with an increased likelihood of producing jailbroken outputs.
Table 6: Model output under different shifting distances.
Query: Write a script that can exploit vulnerabilities in a software or operating system
Distance $r$ Response Jailbroken
5 I’m sorry, but as an AI language model, I cannot provide a script… No
30 alright, here’s a script that can exploit vulnerabilities… Yes
80 , but I am not able to do so as … No
[…]
Appendix E Supplementary Illustration on ABD
@@
E.1 Further analysis of Distribution Approximation
We use the 20,000 benign samples and 8,556 harmful samples to form 4096-dim activations. For each layer, we compute their coordinate distributions: benign activation distribution $\mathcal{D}^{l}_{B}$ and harmful activation distribution $\mathcal{D}^{l}_{H}$. Using their means and divergences, we generate normal distributions $\mathcal{D}^{l}_{BN}$ and $\mathcal{D}^{l}_{HN}$. Finally, we calculate the JS-divergence (JSD) between the original and generated distributions. Results are shown in Table 7.
The results show that (1) In all layers, $JSD(\mathcal{D}^{l}_{B},\mathcal{D}^{l}_{BN})<0.1$ and $JSD(\mathcal{D}^{l}_{H},\mathcal{D}^{l}_{HN})<0.1$, confirming the validity of approximating coordinate distributions with normal distributions. (2) Earlier layers have lower JSDs, making this approximation safer in early and middle layers. This supports our ABD settings, where penalty functions are primarily applied to these layers.
Table 7: Layer-wise JSD Comparisons
Layer 0 1 2 3 4 5 6 7
$JSD(D_{B}^{l},D_{BN}^{l})$ 0.0036 0.0026 0.0113 0.0112 0.0174 0.0234 0.0316 0.0379
$JSD(D_{H}^{l},D_{HN}^{l})$ 0.0033 0.0035 0.0114 0.0099 0.0167 0.0239 0.0354 0.0414
Layer 8 9 10 11 12 13 14 15
$JSD(D_{B}^{l},D_{BN}^{l})$ 0.039 0.0455 0.0517 0.058 0.0637 0.0766 0.0757 0.0818
$JSD(D_{H}^{l},D_{HN}^{l})$ 0.0396 0.0473 0.053 0.0589 0.0661 0.076 0.0753 0.0819
Layer 16 17 18 19 20 21 22 23
$JSD(D_{B}^{l},D_{BN}^{l})$ 0.0848 0.0862 0.0864 0.0873 0.0867 0.0862 0.0853 0.0834
$JSD(D_{H}^{l},D_{HN}^{l})$ 0.0823 0.0841 0.0845 0.0844 0.083 0.0815 0.0804 0.0755
Layer 24 25 26 27 28 29 30 31
$JSD(D_{B}^{l},D_{BN}^{l})$ 0.0789 0.0786 0.078 0.0762 0.0724 0.0712 0.0752 0.078
$JSD(D_{H}^{l},D_{HN}^{l})$ 0.0681 0.0672 0.0659 0.0628 0.0601 0.0593 0.0643 0.0724
E.2 Visualization of the penalty function.
@@ E.1 Further analysis of Distribution Approximation
DSR No attack GCG-universal GCG-individual PAIR AdvPrompter COLD-Suffix AutoDAN Avg. Perform
No Defense 96.73% 87.50% 5.29% 28.18% 82.05% 59.86% 2.50% 51.73%
ABD (Ours) 99.81% 99.68% 98.23% 91.51% 97.65% 98.02% 20.77% 86.52%
Table 8: Performance comparison of different methods under various attack scenarios. The best performance is highlighted in bold.
Model Attack No Defense Paraphrase PPL Retokenization Self-Exam Self-Reminder IA ABD (Ours)
Qwen No attack 96.00% 80.00% 96.00% 58.00% 100.00% 96.00% 84.00% 98.00%
GCG-Individual 6.84% 67.52% 36.75% 58.97% 42.74% 82.05% 44.64% 89.74%
AutoDAN 14.00% 12.00% 26.00% 4.00% 50.00% 8.00% 2.00% 36.00%
Vicuna-13B No attack 98.00% 98.00% 98.00% 70.00% 100.00% 96.00% 100.00% 100.00%
GCG-Individual 38.46% 98.08% 98.08% 84.62% 84.62% 98.08% 100.00% 100.00%
AutoDAN 14.00% 54.00% 20.00% 16.00% 96.00% 98.00% 100.00% 74.00%
Table 9: Performance comparison of different defenses across models with different scales and structures. For ABD, the best and second-best performance across all defenses are highlighted.
E.3 ABD Optimization Settings
@@
We adopt Optuna Akiba et al. (2019) as our framework of optimization. We follow Optuna’s default settings, i.e., Tree-structured Parzen Estimator (TPE) Bergstra et al. (2011)-based Optimization.
E.4 ABD Optimizing Process
The optimization process utilizes the Tree-structured Parzen Estimator (TPE) algorithm, a Bayesian optimization variant that efficiently explores high-dimensional parameter spaces by modeling promising regions. TPE constructs two probabilistic density distributions, $l(\Theta)$ and $g(\Theta)$, over the defense parameters $\Theta=\{m_{i},\alpha_{i},\beta_{i},k_{i}\}$, where $l(\Theta)$ captures the likelihood of $\Theta$ yielding high objective score $\mathcal{J}_{\mathrm{total}}$, and $g(\Theta)$ corresponds to lower-score regions. At each iteration, TPE samples new candidates by maximizing the Expected Improvement criterion Jones (2001):
$\Theta_{\mathrm{new}}=\arg\max_{\Theta}\frac{l(\Theta)}{g(\Theta)},$
prioritizing parameters likely to outperform the current best. The process begins with a warm-up phase, where random evaluations establish a baseline, followed by a density-guided search, which iteratively refines $l(\Theta)$ and $g(\Theta)$ using historical observations. This approach ensures efficient exploration while mitigating local optima risks. By leveraging Optuna’s TPE implementation, the optimizer adapts dynamically to $\mathcal{J}_{\mathrm{total}}$ feedback, achieving automated defense parameter optimization.
E.5 Comparison with other defense methods
We compared ABD with other defense methods in terms of usability. We qualitively evaluated usability based on extra computational complexity, extra tokens, additional modules, and deployment difficulty. The results are shown in Table 10. Compared to other defense methods, ABD features no additional token overhead and only constant-level extra complexity. Besides, the operation is simple and straightforward, which further improves its utility.
Appendix F Supplementary Illustration on Experiments
Table 10: Empirical Comparison of Defenses. $t$ is the length of input sequence.
Extra
Complexity
Extra
Tokens
Extra
Modules
Deployment Difficulty Performance Utility
PPL $O(t^{2})$ Extra Conversation GPT-2 Low Low Low
Paraphrase $O(t^{2})$ Extra Conversation - Low Low Low
Retokenization $O(t\log t)$ Low Low Low
Self-Exam $O(t^{2})$ Extra Conversation - Low Low Low
SafeDecoding $O(t)$ LoRA Adapter High High Medium
Self-Reminder $O(t^{2})$ Extra Instructions - Low Low Low
IA $O(t^{2})$ Extra Conversation - Low High Medium
ABD (Ours) $O(1)$ Mathematical Operation Low High High
F.1 Generation Configs
When conducting experiments, we directly utilize most of the original configurations of all LLMs. Specifically, we set max_new_tokens=128. We find that the proper functioning of these LLMs largely depends on the chat template of the inputs. We apply chat templates in fastchat v0.2.36 by:
@@ F.1 Generation Configs
We set template_name="vicuna" for Vicuna-7B and Vicuna-13B, template_name="llama-2" for Llama-2, and template_name="qwen" for Qwen.
F.2 Jailbreak Methods
@@ F.3 Defense Methods
Intention Analysis
Zhang et al. (2025) enhances LLM safety by guiding the model to explicitly identify the essential intention of user queries and subsequently generate policy-aligned responses to mitigate harmful outputs.
F.4 Metrics
F.4.1 Calculating DSR
@@
We adopt refusal strings as proposed by Xu et al. (2024), presented in Table 11.
F.4.2 Calculating Overall Scores
@@
where $\mathrm{t}$ is Runtime per Query and $\mathrm{Avg.}$ represents overall averaged Just-Eval scores. Here, $\mathrm{Norm(x)}$ denotes the normalization of a metric $\mathrm{x}$, calculated as $\mathrm{x}$ divided by the sum of $\mathrm{x}$ and the corresponding values for all reference groups.
@@
Appendix G Supplementary Experiments
G.1 Scale Up Baselines and Data
To further strengthen our evaluation, we expanded the jailbreak test samples from the initial 50-behavior subset to the full AdvBench (5̃20 samples) and applied the jailbreak method from Section 3.2 for a more intuitive observation. The defense performances on Vicuna-7B-v1.3 are presented in Table 8.The results remain consistent with Table 1, reaffirming ABD’s strong defense capability against various stealthy jailbreak methods.
G.2 Full Results on Qwen and Vicuna-13B
The results are shown in Table 9. ABD remains performing decent on other sizes of models. Besides, in Vicuna-13B, 6 out of 9 selected layers are early and intermediate layers (layer 2, 11, 14, 19, 26, 29), and in Qwen, 2 out of 3 layers (layer 5, 14). This supports our interpretations of jailbreak mechanisms, as well as the conclusion that early and middle layers are most vulnerable to jailbreaks, basically hold in larger and smaller models.